OpenStudy (anonymous):

Use substitution to solve for y in the system of equations. 3x - 2y = 33 4x + y =33

ganeshie8 (ganeshie8):

solve for y from second equation

OpenStudy (anonymous):

5?

ganeshie8 (ganeshie8):

second equation :- 4x + y =33

ganeshie8 (ganeshie8):

to solve y, u need to isolate y ok

OpenStudy (anonymous):

i guess

ganeshie8 (ganeshie8):

4x + y =33 subtract 4x both sides 4x + y =33 -4x -4x

ganeshie8 (ganeshie8):

wat do u get ?

ganeshie8 (ganeshie8):

humm shud be simple :- 4x + y =33 -4x -4x 4x cancels out... leaves u wid :- y = 33-4x

ganeshie8 (ganeshie8):

substitute this y value in first equation :- 3x - 2(33-4x) = 33 see if u can solve x :)

OpenStudy (anonymous):

why don't u simply subtract the equations